What are the values of the components of the 53kN Force directed up to the left at an angle of 25 degrees with respect to x-axis?
1
Expert's answer
2021-08-25T17:23:24-0400

Given:

"F=53\:\rm kN"

"\theta=25^{\circ}"


The components of the force

"F_x=F\cos\theta=53\:\rm kN*\cos 25^{\circ}=48\: kN"

"F_y=F\sin\theta=53\:\rm kN*\sin 25^{\circ}=22\: kN"